solve:- (x+1)(x+2)(x+3)(x+4)+1=0

solve:-


(x+1)(x+2)(x+3)(x+4)+1=0

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sayan

(x+1)(x+2)(x+3)(x+4)+1=0

(x+1)(x+1+1)(x+3)(x+3+1)+1=0

[ (x+1)2 +(x+1) ][ (x+3)2 +(x+3)] +1 =0

(x+1)2(x+3)2 +(x+3)2(x+1) +(x+1)2(x+3) +(x+1)(x+3) +1=0

(x+1)2(x+3)2   +2(x+1)(x+3)(x+2) +(x+2)2 =0

[(x+1)(x+3) +(x+2)]2 =0

 [(x+1)(x+3) +(x+2)]=0

 x2 +5x +5 =0

 now solve for x

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin